OpenStudy (anonymous):

Solve The Equation. Log(x+7)-Logx=3 A.)3.5 B.)142.7 C.)0.0070 D.)0.0707

zepdrix (zepdrix):

Rule of logs:\[\Large\rm \log(a)-\log(b)=\log\left(\frac{a}{b}\right)\]

zepdrix (zepdrix):

Applying this rule gives us:\[\Large\rm \log\left(\frac{x+7}{x}\right)=3\]

zepdrix (zepdrix):

From here, you want to convert this to exponential form. Do you know how to do that?

OpenStudy (anonymous):

Kinda.

zepdrix (zepdrix):

You can `exponentiate` each side, which is a kind of a weird process if you haven't done it before. Let's look at my silly rule for converting from log to exponent form:

OpenStudy (anonymous):

Okkie Dokkie. :)

zepdrix (zepdrix):